How many quarts of water must be added to 9 qt of an 80% antifreeze solution to make a 45% antifreeze solution?
Answer by mananth(16946)

let water required be x qt. 0%
9 qt 80%
Mixture = (x+9) qt 45%
..
9*0.8 = (x+9)*0.45
7.2=0.45x+4.05
0.45x=7.20-4.05
divide by 0.45
x= 3.15/0.45
x= 7 qt of water